OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

this one too @jim_thompson5910 please

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = 2 (-1,5) x1 = -1 and y1 = 5 now lets sub our stuff in... y - 5 = 2(x - (-1) y - 5 = 2(x + 1) y - 5 = 2x + 2 y = 2x + 2 + 5 y = 2x + 7 standard form : y = 2x + 7 -2x + y = 7 (multiply by -1) 2x - y = -7

OpenStudy (texaschic101):

any questions ?

OpenStudy (texaschic101):

The reason why I multiplied by -1 is because we do not want x being negative

OpenStudy (anonymous):

ok @texaschic101 yea i see that but i dont undestand why

OpenStudy (texaschic101):

Ax + By + C is the standard formula. " A" is not supposed to be negative....I am not sure if that is a rule or not, but I try not to let it be negative.

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (anonymous):

can you help me with this one too @texaschic101

OpenStudy (texaschic101):

for this one we need to find the slope by using the slope formula. slope(m) = (y2 - y1) / (x2 - x1) (1,3) x1 = 1 and y1 = 3 (4,4) x2 = 4 and y2 = 4 now lets sub.. m = (4 - 3) / (4 - 1) m = 1/3 now we use this one : y - y1 = m(x - x1) and we can use either of the points I will use (1,3) and slope 1 y - 3 = 1/3(x - 1) y - 3 = 1/3x - 1/3 y = 1/3x - 1/3 + 3 y = 1/3x - 1/3 + 9/3 y = 1/3x + 8/3 standard form : y = 1/3x + 8/3 (multiply by 3 to get rid of fractions) 3y = x + 8 -x + 3y = 8 (multiply by -1) x - 3y = -8

OpenStudy (anonymous):

wait is this called point-slope formula?

OpenStudy (texaschic101):

I used the point slope formula : y - y1 = m(x - x1) and then turned it into standard form

OpenStudy (texaschic101):

I used the point slope because we had a point and we had the slope